Explain what is mearnt by the root mean square velocity of the molecules of a gas.
1
Expert's answer
2021-02-09T10:40:45-0500

By definition, the root mean square velocity (speed) is (https://en.wikipedia.org/wiki/Maxwell–Boltzmann_distribution#Typical_speeds):


"v_{rms} = \left( \int_0^{\infty }v^2f(v)dv \right)"

where

"f(v) = \left(\frac{m}{2 \pi kT}\right)^{3/2}\, e^{ - \frac{mv^2}{2kT}}"

is the Maxwell–Boltzmann distribution. The value of root mean square velocity is:


"v_{rms} = \sqrt{\dfrac{3RT}{M}}"

where "R" is the gas constant, "T" is the absolute temperature of the gas, and "M" is the molar mass of the gas.